Northwestern Memorial Hospital is planning an advertising campaign. It cost $3000 each time a television ad is aired and $2000 each time a radio ad is aired. The administrator wants to air 60 more television ads than radio asds. If the total cost of airing the ads is $580,000, then how many ads of each ype will be aired?
Answer by doukungfoo(195)

Let call the radio ads x
Define tv ads in terms of x
tv ads equals x+60
Given: total cost of both ads is 580000
Given: radio ads cost 2000
Given: tv ads cost 3000
2000(x)+3000(x+60)=580000
2000x+3000x+180000=580000
5000x+180000=580000
5000x=400000
x=80
Answer: 80 radio ads
Find tv ads:
x+60
80+60
140
Answer: 140 tv ads
Check:
tv ads times cost plus radio ads times cost equals 580000
(140*3000)+(80*2000)=580000
420000+160000=580000
580000=580000